OpenStudy (anonymous):

Find the horizontal asymptote of the function (4x-8)/((x-4)(x+1))

OpenStudy (anonymous):

\[\frac{ 4x-8 }{ (x-4)(x+1) }\]

OpenStudy (anonymous):

please help me!

OpenStudy (omnomnom):

o.o idk know this sorry T_T

OpenStudy (anonymous):

What is the long-term behavior? i.e. What happens at x=∞ and x=-∞?

OpenStudy (anonymous):

Here's a quick way to figure it out: First, what are the degrees of the numerator and denominator?

OpenStudy (anonymous):

1=numerator 2=denominator

OpenStudy (anonymous):

So as x approaches infinity, you'll have something like x/x^2 which gets bigger faster, x or x^2? It's going to ∞/∞^2

OpenStudy (anonymous):

y=0

OpenStudy (anonymous):

Why do you say that, @michelle1503 ?

OpenStudy (anonymous):

if the degree of the numerator is smaller than the degree of the denominator, the horizontal asymptote is y=0.... its a rule

OpenStudy (anonymous):

Thanks michelle!

OpenStudy (anonymous):

my pleasure :)

OpenStudy (anonymous):

@Brent0423 do you understand based on what I said why what michelle said is a rule?

OpenStudy (anonymous):

can u help me with (18)/((x-3)(x-3))

OpenStudy (anonymous):

i have to find x and y intercepts and horizontal/vertical asymptotes

OpenStudy (anonymous):

@CliffSedge yes I do!

OpenStudy (anonymous):

do u rather want me to give u all the easy rules to find all those answers for future questions? will make life sooo much easier and ill help u with this one as an example :)

OpenStudy (anonymous):

to find the y int u set x=0 and solve for y so i did and i found y=2 so theres an intercept at y=2 for the x-int i set y=0 and solved for x and i found that theres no x-int

OpenStudy (anonymous):

@michelle1503 that would be great! :)

OpenStudy (anonymous):

u are correct with ur intercepts :)

OpenStudy (anonymous):

horizontal asymptote y=0

OpenStudy (anonymous):

and theres no vertical asymptote

OpenStudy (anonymous):

Oh, there is a vertical asymptote alright. . .

OpenStudy (anonymous):

vertical asymptote x=3

OpenStudy (anonymous):

RULES: * to find y intercepts, make all x in the equation = 0 * to find x intercepts, make y= the NUMERATOR and solve when y =0 * to find a vertical asymptote, make DENOMINATOR =y and solve for y=0 *to find a horizontal asymptote, its gets serioushaha:) there are 3 rules.... ......1....if the deg of numerator is smaller that deg of denominator, HA= y=0 ......2....if the deg of numerator is bigger than deg of denominator, there is no horizontal asymptote.. (you get a slanted asymptote but its fine no need to worry about that now:D) .......3....if deg of numerator is the same as the deg of denominator, then the coefficients of the highest deg are divided by eachother (e.g....4x^2/2x^2... deg are equal therefore i say that 4/2=2.. therefore y=2 is HA.) yes u are correct that HA is y=0 now to find ur VA u need to make ur DENOMINATOR =to 0... sooo it will look like this...\[(x-3)(x-3)=0 \] now solving that will give u.... x=3 as your vertical asymptote

OpenStudy (anonymous):

hope that sheds some light :)

OpenStudy (anonymous):

hahaha yeah u got the VA before i could finish typing haha.. awesome :)

OpenStudy (anonymous):

thanks so much for your help with all of this! :)

OpenStudy (anonymous):

no problemo :)